OpenStudy (anonymous):

Yesterday, the price of a loaf of bread at five supermarkets was $1.54, $1.69, $1.79, $1.57, and $1.71, but today each supermarket raised its price by $0.05. What is today's mean price of a loaf of bread at the five supermarkets? A. $1.66 B. $1.81 C. $1.71 D. $1.76 pretty sure its a tell me if im right

OpenStudy (anonymous):

It's not, because you forgot to add 5 cents to everything

OpenStudy (anonymous):

oh thank you

OpenStudy (simoner):

1.59+1.74+1.84+1.62+1.76=8.55/5=1.71

OpenStudy (anonymous):

got it ty